# Geometric measures of CHSH quantum nonlocality: characterization, quantification, and comparison by distances and operations

### v4

@@ IV.0.2 Relative entropy
$\mathcal{M}_{Re}(\rho)\geq\displaystyle\sum_{i}p_{i}\mathcal{M}_{Re}(\rho_{i}),$ (16)
@@
where $\{\Lambda{i}\}$ is any set of Kraus operators, with weights $p_{i}={\rm Tr}(\Lambda_{i}\rho\Lambda_{i}^{\dagger})$, under the hypothesis that the involved Kraus operators are free, i.e. preserve the local states: $\sum_{i}\Lambda_{i}^{\dagger}\Lambda_{i}=\mathbb{I}$ and $\Lambda_{i}\rho\Lambda_{i}^{\dagger}\in\sigma$ $\forall\rho^{loc}\in\sigma$, where $\sigma$ is the set of local states.
@@
To obtain the last inequality, first of all note that since the relative entropy is invariant under the same unitary or isometry applied to both arguments, we can use the isometry $I=\sum_{i}|i\rangle_{aux}\otimes|i\rangle_{aux^{\prime}}\otimes V_{i}$ for any $V_{i}$ to derive the identity $S(\rho^{(1)}|\rho^{(2)})=S(I\rho^{(1)}I^{\dagger}|I\rho^{(2}I^{\dagger})$, for any couple of states $\rho^{(1)}$ and $\rho^{(2)}$, where $\{|i\rangle_{aux}\}_{i}$ is an orthogonal basis of an auxiliary system. Next, monotonicity of the relative entropy under the partial trace over the auxiliary system $aux$ implies
$S(\rho^{(1)}|\rho^{(2)})\geq\sum_{i}S(\tilde{\rho}_{i}^{(1)}|\tilde{\rho}_{i}^{(2)}),$ (17)
@@
where we use unnormalized states $\tilde{\rho}_{i}^{(1)}=V_{i}\rho^{(1)}V_{i}^{\dagger}$ and $\tilde{\rho}_{i}^{(2)}=V_{i}\rho^{(2)}V_{i}^{\dagger}$. Following the argument derived in Vedral1998 we have
$\sum_{i}S(\rho_{i}^{(1)}|\rho_{i}^{(2)})\geq\sum_{i}p_{i}S\left(\frac{\tilde{\rho}_{i}^{(1)}}{p_{i}}\bigg{|}\frac{\tilde{\rho}_{i}^{(2)}}{q_{i}}\right),$ (18)
@@
where $p_{i}={\rm Tr}\ \tilde{\rho}_{i}^{(1)}$ and $q_{i}={\rm Tr}\ \tilde{\rho}_{i}^{(2)}$. Identifying $\rho\equiv\rho^{(1)}$, assuming that $\rho^{loc}\equiv\rho^{(2)}$ is a local sate and $V_{i}$ are Kraus operators which preserve locality, such that $V_{i}\rho^{loc}V_{i}^{\dagger}/q_{i}$ is still a local state, we conclude that
$\mathcal{M}_{Re}(\rho)\geq\sum_{i}p_{i}\mathcal{M}_{Re}\left(\frac{V_{i}\rho V_{i}^{\dagger}}{p_{i}}\right).$ (19)
IV.1 Distances and induced measures of nonlocality
Here we discuss the properties of the geometric based measures of Bell nonlocality as defined, respectively, by Eq. (5) and Eq. (15). Concerning the geometric measures, we will consider those associated to the Hilbert-Schmidt norm and to some of the most relevant contractive distances. In general, each distance provides a slightly different clue on the distinguishability of quantum states.
IV.1.1 Hilbert-Schmidt distance
The Hilbert-Schmidt distance $D_{HS}(\rho_{1},\rho_{2})$ between two density operators $\rho_{1}$ and $\rho_{2}$ is defined trough the Hilbert-Schmidt norm as follows:
$D_{HS}(\rho_{1},\rho_{2})\,=\,\|\rho_{1}-\rho_{2}\|_{HS}\,=\,\sqrt{{\rm Tr}\left[(\rho_{1}-\rho_{2})(\rho_{1}-\rho_{2})\right]}\,\,.$ (20)
For convenience, we define the corresponding geometric measure of Bell nonlocality $\mathcal{M}_{HS}(\rho)$ through the squared distance:
$\mathcal{M}_{HS}(\rho)\,=\,\min_{\rho^{loc}\in\sigma}D_{HS}^{2}(\rho,\rho^{loc})\,.$ (21)
The Hilbert-Schmidt (HS) distance has no direct operational meaning and in general may not monotonic under all CPTP transformations Bengtsson2017 ; Piani2012 . For instance, it increases under a trivial extension of local subsystems. On the other hand, it provides easily computable bounds on valid contractive distances that are often more difficult to compute.
IV.1.2 Hellinger distance
$D_{He}(\rho_{1},\rho_{2})=\|\sqrt{\rho_{1}}-\sqrt{\rho_{2}}\|_{2}.$ (22)
$D_{He}(\rho_{1},\rho_{2})=\sqrt{2-2\sum_{i}\sqrt{e^{(1)}_{i}e^{(2)}_{i}}}=\sqrt{\displaystyle\sum_{i=1}^{4}\bigg{(}\sqrt{e_{i}^{(1)}}-\sqrt{e_{i}^{(2)}}\bigg{)}^{2}}.$ (23)
The above is the Euclidean norm in the $\sqrt{e_{i}}-$space and coincides with the classical Hellinger distance quantifying the distance between the two classical probability distributions Ref_Hellinger1909 . The corresponding Hellinger measure of Bell nonlocality is thus defined as
$\mathcal{M}_{He}(\rho)=\min_{\rho^{loc}\in\sigma}D_{He}^{2}(\rho,\rho^{loc})\,.$ (24)
IV.1.3 Bures distance
The Bures distance is defined as Bengtsson2017
$D_{B}=\sqrt{2\big{[}1-\sqrt{F(\rho_{1},\rho_{2}})\big{]}}\,,$ (25)
where $F$ is the Uhlmann fidelity given by Uhlmann1976 ; Jozsa1994 ; Nielsen2005
$F(\rho_{1},\rho_{2})=\bigg{(}\text{Tr}\bigg{[}\sqrt{\sqrt{\rho_{1}}\rho_{2}\sqrt{\rho_{1}}}\bigg{]}\bigg{)}^{2}\,.$ (26)
$\mathcal{M}_{B}(\rho)=\min_{\rho^{loc}\in\sigma}D_{B}(\rho,\rho^{loc})\,.$ (27)
IV.1.4 Trace distance
The trace distance is defined as
$D_{Tr}(\rho_{1},\rho_{2})=\frac{1}{2}\|\rho_{1}-\rho_{2}\|_{1}=\frac{1}{2}{\rm Tr}\left[\sqrt{(\rho_{1}-\rho_{2})(\rho_{1}-\rho_{2})}\right]\,,$ (28)
$D_{Tr}(\rho_{1},\rho_{2})=\frac{1}{2}{\rm Tr}|\rho_{1}-\rho_{2}|=\frac{1}{2}\sum_{i}|e_{i}^{1}-e^{2}_{i}|\,.$ (29)
The corresponding trace-distance measure of Bell nonlocality reads
$\mathcal{M}_{Tr}(\rho)=\min_{\rho^{loc}\in\sigma}D_{Tr}(\rho,\rho^{loc}).$ (30)
The trace distance defines the probability of error in the Helstrom measurement which is the optimal measurement to distinguish two quantum states. It is jointly convex and nonincreasing under CPTP maps Bengtsson2017 ; Nielsen2005 . As a consequence, the corresponding measure of Bell nonlocality Eq. (30) satisfies monotonicity under LOSR and monotonicity under mixing (convexity).
V Application to two-qubit states: Formalism and CHSH locality
In this section we specialize to two-qubit states in relation to the CHSH inequality Eq. (3). Before proceeding, we recollect some basic facts and derive some theorems that are necessary in order to apply the general framework outlined in the previous sections to the concrete instance of two-qubit states. An arbitrary two-qubit state $\rho$ can be represented as
@@ V Application to two-qubit states: Formalism and CHSH locality
where $\sigma_{0}=\mathbf{1}$ and the $\sigma_{i}$ ($i=1,2,3$) are the Pauli matrices satisfying ${\rm Tr}\sigma_{i}=0$, ${\rm Tr}\sigma_{i}\sigma_{j}=\delta_{ij}$, and the superposition coefficients read $\alpha_{00}=1$, $\alpha_{0i}=\text{Tr}\big{[}\rho\big{(}\mathbb{I}\otimes\sigma_{i}\big{)}\big{]}$, $\alpha_{i0}=\text{Tr}\big{[}\rho\big{(}\sigma_{i}\otimes\mathbb{I}\big{)}\big{]}$ and $\alpha_{ij}=\text{Tr}\big{[}\rho\big{(}\sigma_{i}\otimes\sigma_{j}\big{)}\big{]}$ for $i=1,2,3$; they can be recast in matrix form as follows:
$\alpha=\begin{bmatrix}1&\vec{\alpha}_{0i}\\ \vec{\alpha}_{i0}&A\end{bmatrix},$ (32)
@@
where all entries are real, $\vec{\alpha}_{0i}$ is a row three-vector of elements $\{\alpha_{0i}\}$, $\vec{\alpha}_{i0}$ is a column three-vector of elements $\{\alpha_{i0}\}$, and $A$ is a $3\times 3$ real matrix. Explicitly:
$\rho=\frac{1}{4}\bigg{(}\mathbf{1}+\sum_{i=1}^{3}\alpha_{0i}\mathbf{1}\otimes\sigma_{i}+\sum_{i=1}^{3}\alpha_{i0}\sigma_{i}\otimes\mathbf{1}+\sum_{i,j=1}^{3}A_{ij}\sigma_{i}\otimes\sigma_{j}\bigg{)}\,.$ (33)
The condition $\text{Tr}[\rho^{2}]\leq 1$ leads to
$\sum_{i}(\alpha_{0i}^{2}+\alpha_{i0}^{2})+\displaystyle\sum_{i,j}\alpha_{ij}^{2}\leq 3\,.$ (34)
It is useful to consider the singular value decomposition of the matrix $A$:

### v5

@@ 3.2 Relative entropy
$\mathcal{M}_{\mathrm{Re}}(\rho)\geq\displaystyle\sum_{i}p_{i}\mathcal{M}_{\mathrm{Re}}(\rho_{i}),$ (22)
@@
where $\{\Lambda{i}\}$ is any set of Kraus operators, with weights $p_{i}={\rm Tr}(\Lambda_{i}\rho\Lambda_{i}^{\dagger})$, under the hypothesis that the involved Kraus operators are free, i.e. preserve the local states: $\sum_{i}\Lambda_{i}^{\dagger}\Lambda_{i}=\mathds{1}$ and $\Lambda_{i}\rho\Lambda_{i}^{\dagger}\in\mathcal{L}$ $\forall\rho^{loc}\in\mathcal{L}$, where $\mathcal{L}$ is the set of local states.
@@
To obtain the last inequality, first of all note that since the relative entropy is invariant under the same unitary or isometry applied to both arguments, we can use the isometry $I=\sum_{i}|i\rangle_{aux}\otimes|i\rangle_{aux^{\prime}}\otimes V_{i}$ for any $V_{i}$ to derive the identity $S(\rho^{(1)}|\rho^{(2)})=S(I\rho^{(1)}I^{\dagger}|I\rho^{(2)}I^{\dagger})$, for any couple of states $\rho^{(1)}$ and $\rho^{(2)}$, where $\{|i\rangle_{aux}\}_{i}$ is an orthogonal basis of an auxiliary system. Next, monotonicity of the relative entropy under the partial trace over the auxiliary system auxaux implies
$S(\rho^{(1)}||\rho^{(2)})\geq\sum_{i}S(\tilde{\rho}_{i}^{(1)}||\tilde{\rho}_{i}^{(2)}),$ (23)
@@
where we use unnormalized states $\tilde{\rho}_{i}^{(1)}=V_{i}\rho^{(1)}V_{i}^{\dagger}$ and $\tilde{\rho}_{i}^{(2)}=V_{i}\rho^{(2)}V_{i}^{\dagger}$. Following the argument derived in [44] we have
$\sum_{i}S(\rho_{i}^{(1)}||\rho_{i}^{(2)})\geq\sum_{i}p_{i}S\left(\frac{\tilde{\rho}_{i}^{(1)}}{p_{i}}\bigg|\frac{\tilde{\rho}_{i}^{(2)}}{q_{i}}\right),$ (24)
@@
where $p_{i}={\rm Tr}\ \tilde{\rho}_{i}^{(1)}$ and $q_{i}={\rm Tr}\ \tilde{\rho}_{i}^{(2)}$. Identifying $\rho\equiv\rho^{(1)}$, assuming that $\rho^{loc}\equiv\rho^{(2)}$ is a local state and $V_{i}$ are Kraus operators which preserve locality, such that $V_{i}\rho^{loc}V_{i}^{\dagger}/q_{i}$ is still a local state, we conclude that
$\mathcal{M}_{\mathrm{Re}}(\rho)\geq\sum_{i}p_{i}\mathcal{M}_{\mathrm{Re}}\left(\frac{V_{i}\rho V_{i}^{\dagger}}{p_{i}}\right).$ (25)
4 Application to two-qubit states
An arbitrary two-qubit state $\rho$ can be represented as
@@ 4 Application to two-qubit states
where $\sigma_{0}=\mathds{1}_{2}$ and $\sigma_{i}$ ($i=1,2,3$) denote the Pauli matrices, satisfying ${\rm Tr}[\sigma_{i}]=0$ and ${\rm Tr}[\sigma_{i}\sigma_{j}]=2\delta_{ij}$. The expansion coefficients are given by $\alpha_{00}=1$, $\alpha_{0i}={\rm Tr}\big[\rho\,(\mathds{1}_{2}\otimes\sigma_{i})\big]$, $\alpha_{i0}={\rm Tr}\big[\rho\,(\sigma_{i}\otimes\mathds{1}_{2})\big]$, and $\alpha_{ij}={\rm Tr}\big[\rho\,(\sigma_{i}\otimes\sigma_{j})\big]$ for $i=1,2,3$. These coefficients can be arranged in matrix form as
$\alpha=\left[\begin{array}[]{cc}1&\vec{\alpha}_{0i}\\ \vec{\alpha}_{i0}&A\end{array}\right],$ (27)
@@
where all entries are real, $\vec{\alpha}_{0i}$ is a row vector with components $\{\alpha_{0i}\}$, $\vec{\alpha}_{i0}$ is a column vector with components $\{\alpha_{i0}\}$, and $A$ is a $3\times 3$ real matrix. As first result, we prove the following:
Result 1. If $\rho$ is a local state with given $(\alpha_{0i},\alpha_{i0},A)$ then the state $\rho^{-}$ given by $(-\alpha_{0i},-\alpha_{i0},A)$ is also local.
Proof. It suffices to consider a local transformation that for $i=1,2,3$ changes $\sigma_{i}$ into $\sigma_{i}$. $\square$
Result 2. For any local state $\rho$ with given $(\alpha_{0i},\alpha_{i0},A)$ there exists a local state $\rho^{\prime}$ with the same $A$ but with maximally mixed reduced states, i.e., with $\alpha_{0i}=\alpha_{i0}=0$
Proof. Consider the convex combination
$\rho^{\prime}=\frac{1}{2}\rho+\frac{1}{2}\rho^{-}.$ (28)
Indeed, for such a combination we have
$\alpha=\left[\begin{array}[]{cc}\alpha_{00}&0\\ 0&A\end{array}\right].$ (29)
Now, convex combinations of local states are local and this ends the proof. $\square$
Two particularly relevant families of states are the Bell-diagonal states, i.e., convex mixtures of the four Bell states (see A for details):
[…]
and the subclass known as Werner states, which are convex combinations of the maximally mixed state and the singlet Bell state. They are parametrized by a single parameter $\omega$ and were introduced as the first example of states that are entangled yet do not necessarily violate the CHSH inequality [10].
For this class of states we prove the following two propositions, holding for every scenario in which a set of Bell inequalities (just “Bell inequality” in the following) defines the convex set $\mathcal{L}$ of local states in that scenario:
Proposition 1. For any Bell inequality, any jointly convex and unitarily invariant functional $D$ of pairs of two-qubit density matrices
$\min_{\rho^{loc}\in\mathcal{L}}D(\rho_{w},\rho^{loc})=\min_{\rho^{loc}_{w}\in\mathcal{L}}D(\rho_{w},\rho^{loc}_{w})\,$, (31)
where $\rho_{w}$ is a Werner state, $\rho^{loc}$ is a local state, and $\rho^{loc}_{w}$ is a local Werner state.
Proof. The proof of Proposition 1 is shown in section 6 for isotropic states and generalized Werner states in higher dimensions. Here we just sketch the relevant steps. Starting from Eq. (4), we can apply the same $U\otimes U$ to both arguments of $D$ and make use of the unitarily invariance of $D$; then, we bound this quantity from below by twirling both arguments and using jointly convexity; finally we recognize that twirling projects $\rho_{w}$ itself and $\rho^{loc}$ into another (local) Werner state $\rho^{loc}_{w}$. So $D(\rho_{w},\rho^{loc}_{w})\leq D(\rho_{w},\rho^{loc})$, hence the closest state can be taken to be a Werner. $\square$
Proposition 2. For any Bell inequality, any jointly convex and unitarily invariant functional $D$ of pairs of two-qubit density matrices, the closest local state to a Bell-diagonal state is Bell-diagonal:
$\min_{\rho^{loc}\in\mathcal{L}}D(\rho_{\mathrm{BD}},\rho^{loc})=\min_{\rho_{\mathrm{BD}}^{loc}\in\mathcal{L}}D(\rho_{\mathrm{BD}},\rho_{\mathrm{BD}}^{loc})\,$, (32)
where $\rho_{\mathrm{BD}}$ denotes a Bell-diagonal state, $\rho^{loc}$ a local state, and $\rho_{\mathrm{BD}}^{loc}$ a local Bell-diagonal state.
Proof.
The symmetry class that we consider contains the simultaneous local $\pi$-rotations of qubits around the three Pauli axes. Rotation $R_{1}$ around $\sigma_{1}$ changes $\sigma_{2}$ and $\sigma_{3}$ to $\sigma_{2}$ and $\sigma_{3}$, respectively. Rotations $R_{2}$ and $R_{3}$ operate analogously. These transformations are local and unitary, therefore, they cannot change the locality of a quantum state. Moreover, these transformations preserve Bell-diagonal states. Indeed, let us consider the general state representation Eq. (26). Under rotation $R_{1}$, we have the following transformation:
$\hskip-28.45274pt\alpha=\left(\begin{array}[]{cccc}1&a_{01}&a_{02}&a_{03}\\ a_{10}&a_{11}&a_{12}&a_{13}\\ a_{20}&a_{21}&a_{22}&a_{23}\\ a_{30}&a_{31}&a_{32}&a_{33}\end{array}\right)\,\,\longrightarrow\,\,\alpha_{1}=\left(\begin{array}[]{cccc}1&a_{01}&-a_{02}&-a_{03}\\ a_{10}&a_{11}&-a_{12}&-a_{13}\\ -a_{20}&-a_{21}&a_{22}&a_{23}\\ -a_{30}&-a_{31}&a_{32}&a_{33}\end{array}\right).$ (33)
$\hskip-28.45274pt\alpha_{1}^{\prime}=\left(\begin{array}[]{cccc}1&a_{01}&0&0\\ a_{10}&a_{11}&0&0\\ 0&0&a_{22}&a_{23}\\ 0&0&a_{32}&a_{33}\end{array}\right).$ (34)
Applying rotation $R_{2}$ to $\rho^{\prime}$ and averaging the result with $\rho^{\prime}$ leads to the state with diagonal representation Eq. (30), i.e., a Bell-diagonal state. The total transformation that projects a given state to a Bell-diagonal state takes the compact form
$\hskip-28.45274pt\rho\,\,\longrightarrow\,\,\frac{1}{2}\left(\frac{1}{2}(\rho+R_{1}(\rho))+R_{2}\left(\frac{1}{2}(\rho+R_{1}(\rho))\right)\right).$ (35)
This transformation obviously preserves the Bell-diagonal states, does not change the locality, and is a composition of joint local unitaries and averaging.
The proof of Proposition 2 is thus immediately completed by straightforwardly adapting the reasoning followed in the proof of Proposition 3. $\square$
Note that the previous result is useful for quantifying nonlocality of any two-qubit system with maximally-mixed subsystems, since those states can be transformed in Bell-diagonal states by means of local unitary transformations which do not change the nonlocality of the state.
Indeed, if one considers the singular value decomposition of the matrix $A$: